# What to do if language models disagree? Black-box model ensembling for textual and visual question answering

### v1

@@ 4.1 Datasets
To demonstrate the data efficiency of our approach, we subsampled four publicly available benchmark datasets. This resulted in four Mini datasets, amounting to 1% of the TQA datasets’ and 10% of the VQA datasets’ original size. Table 2 presents the details of these datasets.
TQA datasets. We generated two Mini datasets, Mini-SDv2 and Mini-NQ, consisting of 1,000 randomly sampled instances from SQuAD-V2 (Rajpurkar et al., 2018) and NQ-Open (Kwiatkowski et al., 2019) train splits, respectively. For Mini-NQ, we followed (Fisch et al., 2019) to use long answers as the context, and short answers as the ground truth answers. The 1,000 samples are divided into train and validation data using an 8:2 ratio, while the trained models are tested on the dev data of the original datasets due to the unavailability of original test data. We use the setup proposed in (Laskar et al., 2023) to generate the answers from LLMs. Concretely, this setting relies on prompts generated by PromptSource (Bach et al., 2022), which we apply to our Mini-SDv2 dataset (unavailable for Mini-NQ). The prompts are in two forms depending on the context: (1) “What is the answer? [Context]; [Question]; If you can’t find the answer, please respond ‘unanswerable’. Answer:”; (2) “Answer the question depending on the context. [Context]; [Question]; If you can’t find the answer, please respond ‘unanswerable’. Answer:”. Differently, Mini-NQ does not contain unanswerable questions and thus we use the prompt “Answer the question depending on the context without explanation. [Context]; [Question]; Answer:”. Note that the answers of LLMs can be greatly influenced by some factors such as the use of different prompts or temperatures. However, our study does not focus on prompt engineering but rather on selecting the optimal base model to generate an answer. We will publicly release our prompts as well as the answers from LLMs for reproducibility.
VQA datasets. Our results (Figure 2) reveal that VQA tasks demand a greater quantity of training samples compared to TQA tasks. Therefore, we constructed Mini-GQA and Mini-Viz datasets using a larger fraction (the dev data) of GQA (Hudson & Manning, 2019) and VizWiz (Gurari et al., 2018) datasets compared to TQA datasets. The resulting Mini-GQA and Mini-Viz were divided into train and validation subsets using 8:2 ratio, while the test subset remained the same as in the original datasets.
Mini Dataset | Source Dataset | Num. | % Dataset |
@@
Mini-SDv2 validation | SQuAD-V2 train | 200 | 0.14 |
@@
Mini-NQ validation | NQ-Open train | 200 | 0.22 |
@@ 4.2 Base Models
We experiment with ensembling GPT-3.5-turbo-0613 (ChatGPT), LLaMA-2-70b-chat (hereinafter referred to as “LLaMA”) (Touvron et al., 2023) and GPT-3.5 text-davinci-003 (hereinafter referred to as “Davinci”) to generate answers for TQA tasks.666GPT3.5 text-davinci-003 is deprecated after our experiments, but this fact does not influence the effectiveness of our method. To tackle VQA tasks, we employ three VQA models (VLMo (Bao et al., 2022), ALBEF (Li et al., 2021a) and BLIP (Li et al., 2022)), which are pre-trained on VQA v2 dataset (Antol et al., 2015). Note that we use the publicly accessible VQA models to save experimental costs, but these models are assumed to be black-box. This means that they can only provide predictions without any logits or confidence scores, which is the restriction assumed for the purpose of our study.
@@
4.3 Baselines
Weighted Voting (WV). We adopt a strategy similar to Schick et al. (2020), where the model accuracy of the train data before training is used as the weight for average weighting. In our case, we use the corresponding accuracy of the base models as the weight for voting.
PageRank (Brin & Page, 2012). We adapt PageRank as a baseline to determine the most suitable answer in a graph where all the answers to one question are connected by their BLEURT (Sellam et al., 2020) similarities.
Overall Local Accuracy (OLA) (Woods et al., 1997). Following (Cruz et al., 2018), we use the k-nearest neighbors algorithm to divide the input space (representations of prompts for TQA, representations of images and questions for VQA) of training data into 7 regions. The overall local accuracy of each base model in different regions is computed as its region competence. The model presenting the highest competence level is selected to predict the label of inputs that fall in the same region.
PairRanker and LLM-Blender (Jiang et al., 2023). Both methods were originally designed for text generation tasks (e.g., machine translation and speech recognition). PairRanker model (DeBERTa (He et al., 2023), 400M parameters) is trained to rank a pair of candidate predictions from two LLMs using multiple optimizing objectives (i.e., log probabilities, BART score, BERTScore, etc). A bubble sort is applied to get the top k (we use k=2) predictions among multiple pairwise comparison results to feed to a GENFUSER model (Flan-T5-XL (Chung et al., 2022), 3B parameters) to generate the final fused prediction. LLM-Blender is a composite of the PairRanker and the GENFUSER model. We use a pre-trained (0-shot) PairRanker and LLM-Blender models which have been trained over massive data (105k) including TQA data as baselines to test on our data.
4.4 Experimental Setup
Evaluation Metric. LLMs tend to generate contextual answers that lead to lower scores in the exact match (EM). Therefore, we mainly use the (per-answer) token-level F1-score from the official evaluation guidance of the datasets as the main evaluation metric for TQA performance. Our results differ from the ones reported in (Laskar et al., 2023; Kocoń et al., 2023) because we do not apply any post-processing, human evaluation or output constraints for the generated answers.
@@ 4.4 Experimental Setup
Setup. We fixed the batch size to the upper limit of the server capacity, while the learning rates and epochs are selected after a grid search on a set of values (learning rates: {e3, 5e4, e4, 5e5, e5, 5e6, e6}, epochs: {3, 5, 10, 15, 20}). Models for TQA are trained for 5 epochs using a learning rate of and batch size of 4. Models for VQA use the same learning rate but a batch size of 16 for 20 epochs. We spent 74 and 290 seconds training 1 epoch on 1,000 samples for TQA and 4,319 samples for VQA respectively. The training was performed on 1 GPU with 16GB memory of a DGX1 server ((Pascal) Tesla P100).
Textual Question Answering | Visual Question Answering | ||||||
Model | Mini-SDv2 | Mini-NQ | Model | Mini-GQA | Mini-Viz | ||
EM | F1 | EM | F1 | ACC | ACC | ||
LLaMA-2-70b-chat | 0.24 | 11.34 | 28.07 | 46.47 | ALBEF | 50.60 | 21.28 |
text-davinci-003 | 52.37 | 58.44 | 52.24 | 69.44 | BLIP | 48.08 | 20.80 |
ChatGPT | 30.89 | 44.95 | 57.53 | 71.54 | VLMo | 48.21 | 19.77 |
Oracle | 58.61 | 66.20 | 64.02 | 79.21 | Oracle | 65.03 | - |
MV | 26.95 | 37.75 | 46.07 | 62.43 | MV | 51.05 | 21.47 |
WV | 52.37 | 58.44 | 57.53 | 71.54 | WV | 52.10 | 19.43 |
PageRank | 25.39 | 37.31 | 51.76 | 68.53 | PageRank | 51.47 | 21.66 |
OLA | 47.90 | 55.59 | 54.70 | 70.05 | OLA | 48.65 | 20.32 |
(0-shot) PairRanker | 7.28 | 19.63 | 35.30 | 53.05 | (0-shot) PairRanker | 47.69 | 20.74 |
(0-shot) LLM-Blender | 4.90 | 21.20 | 1.03 | 25.06 | (0-shot) LLM-Blender | 0.0 | 0.0 |
FT-TT | 46.80 | 47.68 | 36.52 | 40.60 | FT-MT | 50.48 | 51.76 |
InfoSel-TT (ours) | 57.74 | 63.63 | 58.45 | 73.37 | InfoSel-MT (ours) | 55.16 | 23.16 |
InfoSel∗-TT (ours) | 49.09 | 49.85 | 48.16 | 53.70 | InfoSel∗-MT (ours) | 52.54 | 52.91 |
4.5 Performance Comparison
@@
In this section, we analyze the performance of our method, taking into account its distinctive characteristics as described in Table 1. Concretely, we focus on comparing our models in terms of task-specific performance, data efficiency, lightweight design, and multimodal capabilities.
@@
Task-specific Performance. Table 3 demonstrates the task-specific performance of InfoSel, base models and baselines on textual and visual QA datasets. For TQA, we observe that LLaMA underperforms other base models. Upon closer examination, we found that LLaMA generates longer explanation text which, although often accurate, decreases the EM and F1-score values. Conversely, a more consistent performance of base models is observed for VQA. All the models demonstrate superior performance on Mini-NQ compared to Mini-SDv2. This is because Mini-SDv2 contains 50% of unanswerable questions, written adversarially and specifically designed to be challenging for the QA task (Rajpurkar et al., 2018). Similarly, Mini-Viz contains 28% of unanswerable questions, and the label “unanswerable” has never been seen by base models. Consequently, this lack of exposure leads to significantly lower performance scores. The Oracle scores of the base model indicate the performance of an ideal ensemble method.
@@
Baseline methods such as WV, PageRank and OLA achieve only marginal improvements compared to base models (+1.5%) on VQA datasets. These results highlight the limitations of these methods when applied to task-specific datasets (see also Table 1). Although 0-shot PairRanker and LLM-Blender models have been trained on massive amounts of data (including question answering data), they exhibit very low performance. LLM-Blender tends to generate longer answers compared to PairRanker which leads to a lower score especially when evaluated in the exact-match settings (EM, ACC).
InfoSel-TT achieves 5.19% (63.63-58.44) improvement compared to individual base models, and reaches 96.12% (63.63/66.20) of the Oracle on Mini-SDv2. Similarly, the corresponding improvement in Mini-NQ performance is 1.83%, reaching 93.06% of the performance achieved by the Oracle. In contrast, FT-TT, despite its superior performance over two base models on Mini-SDv2, underperforms InfoSel-TT by more than 15% due to the small-size training data (refer to Figure 2). We hypothesize that this poor performance of FT-TT models has a direct impact on the performance of InfoSel∗, which did not bring an additional improvement upon InfoSel. Furthermore, Table 3 showcases that InfoSel-MT achieves an improvement of 4.56% in accuracy score compared to the base models (55.16-50.60) on Mini-GQA, reaching 84.81% (55.16/65.03) of the Oracle performance. Furthermore, FT-MT improves 30.48% (51.76-21.28) accuracy on Mini-Viz due to additional new labels (e.g., “unanswerable”) introduced during fine-tuning. The base models in InfoSel-MT lacked exposure to these labels, leading to lower scores compared to FT-MT. Finally, the superior performance of InfoSel∗-MT model on Mini-Viz dataset demonstrates the effectiveness of the proposed blending approach, which improves 31.63% (52.91-21.28) accuracy upon the InfoSel-MT model. Thus, we conclude that the obtained results indicate that InfoSel not only improves the task-specific performance of the base models over four datasets, but also outperforms all the baselines.
@@
Data Efficiency. The experimental results shown in Figure 2 demonstrate the data efficiency of our method by evaluating the model’s performance across varying training data sizes. We observe that InfoSel-TT achieves a higher F1-score compared to the base models when trained on as little as 10 samples from Mini-SDv2. This result has been further verified with the mean F1-score of 10 test results using different seed variations for sampling the training data. Conversely, the number of training of samples needed to surpass the performance of base models is higher for VQA datasets: 5% (6,603 samples) for Mini-GQA and 20% (864 samples) for Mini-Viz. We hypothesize that this is due to the inherent complexity of the VQA task.
@@
Additionally, we find that a larger training data size benefits FT-TT more than InfoSel-TT and OLA. For example, the F1-score of FT-TT increases 200% and 500% from 10 to 10,000 training samples on Mini-SDv2 and Mini-NQ respectively, while InfoSel-TT only increases only 3% and 4%. However, FT-TT still underperforms the best base model, which suggests that fine-tuning a small-size model requires larger training data for getting a comparable performance with LLMs or InfoSel. Finally, we observe that a larger training data size does not necessarily lead to improved performance for the fine-tuned FT-TT model (e.g., when increasing from 80% to 100% the training data size on Mini-GQA). In contrast, OLA does not benefit as much as InfoSel and FT from a larger size of training data, only outperforming InfoSel-TT on Mini-NQ when 10 and 20 training samples are used.
Lightweight Model. Table 4 reports the parameter size of the base models and their ensemble models (InfoSel-TT and InfoSel-MT). InfoSel provides an efficient method for ensembling large LLMs such as ChatGPT (175B parameters) using only 110M trainable parameters. Even though only 37% ((182M-115M)/182M) trainable parameters are saved for the VQA task, we still demonstrate that InfoSel can effectively enhance the task-specific performance of small-size black-box VQA models, offering a lightweight solution.
LLMs | VQA Models | ||
Model | #Param | Model | #Param |
LLaMA-2-70b-chat | 70B | ALBEF | 290M |
text-davinci-003 | 175B | BLIP | 361M |
ChatGPT | 175B | VLMo | 182M |
InfoSel-TT | 110M | InfoSel-MT | 115M |
@@
Multimodal Data. InfoSel is able to utilize multimodal data (image and text) for VQA tasks, and thus outperform the latest text-exclusive LLM ensemble methods (PairRanker and LLM-Blender) as evidenced in Table 3. In contrast, LLM-Blender cannot process image features, thereby lacking crucial information in the multimodal setting, leading to an accuracy of 0.0 on VQA datasets.777The most frequent answer of LLM-Blender on VQA datasets is “I’m sorry, I don’t have enough context to answer that question.” Further insights into the significance of modality information are elaborated in Section 4.6 and Table 5.

### v2

@@ 4.1 Datasets
To demonstrate the data efficiency of our approach, we subsampled four publicly available benchmark datasets. his process resulted in the creation of four Mini datasets. The training set comprises less than 1% of the original size of the TQA datasets and about 10% of the original size of the original VQA datasets. Table 2 presents the statistics of these datasets.
Mini Dataset | Source Dataset | Num. | % |
@@
Mini-SDv2 val | SQuAD-V2 train | 200 | 0.14 |
@@
Mini-NQ val | NQ-Open train | 200 | 0.22 |
@@
4.1.1 TQA Datasets
SQuAD-V2 [31] stands for Stanford Question Answering Dataset 2.0, a dataset designed for the task of question answering. It is an extension of the original SQuAD dataset by including over 50,000 unanswerable questions written adversarially by crowdworkers. The dataset is widely used in natural language understanding research.
NQ-Open [22] is derived from Natural Questions and serves as an open-domain question-answering evaluation. The entirety of the questions can be addressed using the information found in the English Wikipedia. It was created for research purposes.
We generated two Mini datasets, Mini-SDv2 and Mini-NQ, consisting of 1,000 randomly sampled instances from SQuAD-V2 [31] and NQ-Open [22] train splits, respectively.
For Mini-NQ, we followed Fisch et al. [15] to use long answers as the context, and short answers as the ground truth answers. The 1,000 samples are divided into train and validation data using an 8:2 ratio, while the trained models are tested on the dev data of the original datasets due to the unavailability of original test data. We use the setup proposed in [23] to generate prompts with PromptSource [4] for the Mini-SDv2 dataset. These prompts integrated with context and questions from the dataset are used to query answers from LLMs. The prompts for Mini-NQ, unavailable in PromptSource, are adapted from the prompts of Mini-SDv2. The details of these prompts are shown in Table 3.
Dataset | Sample Prompts |
[…]
4.1.2 VQA Datasets
GQA [19] is a large-scale dataset for visual reasoning and compositional question answering research. The dataset contains over 113k images collected from a diverse set of sources and over 22 million questions. Only one ground-truth answer is provided for each image-question pair.
VizWiz [17] is a benchmark dataset for visual question answering. It includes 31K images, 250K questions, and answers collected through a mobile app for visually impaired users. 10 ground-truth answers are provided for each image-question pair.
Mini-GQA and Mini-Viz are sampled from GQA [19] and VizWiz [17]. As the training data of the VQA models are publicly accessible in contrast to LLMs, we compare the label differences of the pre-trained dataset (VQA v2 [2]) of VQA base models with task-specific datasets (GQA, VizWiz) for ensemble training. Figure 2 shows the top 7 most frequent answers and their percentages in GQA, VQA v2 and VizWiz. Four answers in GQA do not appear in the top list of VQA v2 and three for VizWiz (including the top frequent answer "unanswerable"). We also sampled 3k of the most frequent answers from each dataset and calculated their percentage of overlap, which is reported on the intersection in the figure. GQA and VizWiz have 32.9 % and 21.6% of overlap with VQA v2 respectively, showcasing significant differences between the pre-trained dataset and task-specific datasets.
Our results (Figure 3) reveal that VQA tasks demand a greater quantity of training samples compared to TQA tasks. Therefore, we constructed the train and validation sets of Mini-GQA and Mini-Viz datasets using a larger fraction (the dev data) of GQA [19] and VizWiz [17] datasets. Similar to the TQA datasets, the train and validation split ratio is 8:2 ratio.
4.2 Baselines
We use Majority Voting (MV), Weighted Voting (WV) [35], PageRank [7], Overall Local Accuracy (OLA) [39], PairRanker and LLM-Blender [20] as baselines.
Weighted Voting (WV). We adopt a strategy similar to Schick and Schütze [35], where the model accuracy of the train data before training is used as the weight for average weighting. In our case, we use the corresponding accuracy of the base models as the weight for voting.
PageRank [7]. We adapt PageRank as a baseline to determine the most suitable answer in a graph where all the answers to one question are connected by their BLEURT [36] similarities.
Overall Local Accuracy (OLA) [39]. Following [12], we employ the k-nearest neighbors algorithm to divide the input space (comprising representations of prompts for TQA, of images and of questions for VQA) of training data into 7 regions. The overall local accuracy of each base model in different regions is computed as its region competence. The model presenting the highest competence level is selected to predict the label of inputs that fall in the same region.
PairRanker [20]. The PairRanker model is trained to rank a pair of candidate predictions from two LLMs using multiple optimizing objectives (i.e., EM, F1). In the end, the pairwise rank probability produced by the PairRanker model is used to select the top 1 answer as the final prediction. We train PairRanker using the same textual transformer backbone as InfoSel for comparison.
LLM-Blender [20]. LLM-Blender is the extension of PairRanker, which feed the top k (we set k to 2 in the experiments) predictions from the PairRanker model to a GENFUSER model (Flan-T5-XL, as described in [11], which has 3 billion parameters) to generate the final fused prediction. We use pre-trained (0-shot) GENFUSER model from the original paper which have been trained over massive data (105k) including TQA data to test on our data.
@@ 4.3 Base Models
We experiment with ensembling GPT-3.5-turbo-0613 (ChatGPT), LLaMA-2-70b-chat (hereinafter referred to as “LLaMA”) [37] and GPT-3.5 text-davinci-003 (hereinafter referred to as “Davinci”) to generate answers for TQA tasks. To tackle VQA tasks, we employ three VQA models (VLMo [5], ALBEF [25] and BLIP [24]), which are pre-trained on VQA v2 dataset [2]. Note that we use publicly accessible VQA models to save experimental costs, but these models are assumed to be black-box.
Model | #Param. | |
TQA | LLaMA-2-70b-chat | 70B |
text-davinci-003 | 175B | |
ChatGPT | 175B | |
FT-BERT | 110M | |
PairRanker | 110M | |
InfoSel-TT | 110M | |
InfoSel∗-TT | 110M | |
VQA | ALBEF | 290M |
BLIP | 361M | |
VLMo | 182M | |
FT-MT | 115M | |
PairRanker | 110M | |
InfoSel-MT | 115M | |
InfoSel∗-MT | 115M |
ChatGPT is a large language model with 175B parameters, it allows you to have human-like conversations and much more with the chatbot. ChatGPT can generate context-based responses to user prompts (questions). However, this model is currently only accessible by cloud APIs.
GPT 3.5 text-davinci-003 is similar to ChatGPT but designed specifically for instruction-following tasks which enables it to respond concisely and more accurately. This model was deprecated for public access after our experiments, but this did not influence the effectiveness of our method.
LLaMA-2-70b-chat [37] is a large language model with 70B parameters. It is fine-tuned on LLaMA 2 with publicly available instruction datasets and over 1 million human annotations, while Llama 2 models are trained on 2 trillion tokens from publicly available online data sources. LLaMA 2 models are currently publicly accessible.
ALBEF [25]444https://github.com/salesforce/ALBEF first encodes the image and text with an image encoder (visual transformer [14]) and a text encoder respectively. Then a multimodal encoder is used to fuse the image features with the text features through cross-modal attention. The V&L representation is trained with objectives of image-text contrastive learning, masked language modeling and image-text matching. Unlike U-VisualBERT, ALBEF uses a 6-layer transformer decoder to generate answers for VQA tasks.
BLIP [24]555https://github.com/salesforce/BLIP utilizes a visual transformer as the image encoder, and a multi-task model (multimodal mixture of encoder-decoder) as a unified model with both understanding and generation capabilities. The model is jointly pre-trained with three vision-language objectives: image-text contrastive learning, image-text matching, and image-conditioned language modeling. Similarly to ALBEF, this method treats the VQA task as an answer generation task.
VLMo [5]666https://github.com/microsoft/unilm/tree/master/vlmo is a unified vision-language pre-training method with Mixture-of-Modality-Experts. VLMO leverages large-scale image and text data to learn joint representations of vision and language. It employs a mixture model to capture diverse interactions between visual and textual information, achieving state-of-the-art performance on various vision-language tasks.
Table 4 reports the parameter size of the base models and their ensemble models (InfoSel-TT and InfoSel-MT). InfoSel provides an efficient method for ensembling large LLMs such as ChatGPT (175B parameters) using only 110M trainable parameters. Even though only 37% ((182M-115M)/182M) trainable parameters are saved for the VQA task, we still demonstrate that InfoSel can effectively enhance the task-specific performance of small-size black-box VQA models, offering a lightweight solution.
4.4 Evaluation Metric
LLMs tend to generate contextual answers that lead to lower scores in the exact match (EM). Therefore, we mainly use the (per-answer) token-level F1-score from the official evaluation guidance of the datasets as the main evaluation metric for TQA performance. Our results differ from the ones reported in [23, 21] because we do not apply any post-processing, human evaluation or output constraints on the generated answers.
@@ 4.4 Evaluation Metric
Model/Method | Input Info. | Mini-SDv2 | Mini-NQ | ||||
EM | F1 | EM | F1 | ||||
LLaMA-2-70b-chat | ✗ | 0.24 | 11.34 | 28.07 | 46.47 | ||
Base | text-davinci-003 | ✗ | 52.37 | 58.44 | 52.24 | 69.44 | |
Models | ChatGPT | ✗ | 30.89 | 44.95 | 57.53 | 71.54 | |
Oracle | - | - | 58.61 | 66.20 | 64.02 | 79.21 | |
MV | ✗ | 26.95 | 37.75 | 46.07 | 62.43 | ||
WV | ✗ | 52.37 | 58.44 | 57.53 | 71.54 | ||
PageRank [7] | ✗ | 25.39 | 37.31 | 51.76 | 68.53 | ||
Baselines | FT-BERT [41] | 46.80 | 47.68 | 36.52 | 40.60 | ||
OLA [39] | 47.90 | 55.59 | 54.70 | 70.05 | |||
PairRanker [20] | 57.28 | 63.33 | 57.96 | 72.21 | |||
(0-shot) LLM-Blender [20] | 4.90 | 21.20 | 1.03 | 25.06 | |||
Ours | ✗ | 57.50 | 63.60 | 59.07 | 73.46 | ||
InfoSel-TT | ✗ | 57.51 | 63.58 | 58.30 | 72.93 | ||
57.74 | 63.71 | 58.45 | 73.37 | ||||
InfoSel∗-TT | 49.09 | 49.85 | 48.16 | 53.70 |
Model/Method | Input Info. | Mini-GQA | Mini-Viz | |||
ACC | ||||||
ALBEF | ✗ | 50.60 | 21.28 | |||
Base | BLIP | ✗ | 48.08 | 20.80 | ||
Models | VLMo | ✗ | 48.21 | 19.77 | ||
Oracle | - | - | ✗ | 65.03 | - | |
MV | ✗ | ✗ | 51.05 | 21.47 | ||
WV | ✗ | ✗ | 52.10 | 19.43 | ||
PageRank [7] | ✗ | ✗ | 51.47 | 21.66 | ||
Baselines | FT-MT [41] | 50.48 | 51.76 | |||
OLA [39] | 48.65 | 20.32 | ||||
PairRank [20] | ✗ | 52.05 | 22.42 | |||
(0-shot) LLM-Blender [20] | ✗ | 0.0 | 0.0 | |||
Ours | InfoSel-MT | ✗ | ✗ | 50.56 | 20.79 | |
✗ | ✗ | 51.11 | 21.21 | |||
✗ | ✗ | 51.69 | 22.83 | |||
✗ | 50.83 | 20.06 | ||||
✗ | 52.38 | 22.66 | ||||
✗ | 54.76 | 22.89 | ||||
55.16 | 23.16 | |||||
InfoSel∗-MT | 52.54 | 52.91 |
4.5 Experiment Setup
@@
We fixed the batch size to the upper limit of the server capacity, while the learning rates and epochs are selected after a grid search on a set of values (learning rates: {e3, 5e4, e4, 5e5, e5, 5e6, e6}, epochs: {3, 5, 10, 15, 20}). Models for TQA are trained for 5 epochs using a learning rate of and batch size of 4. Models for VQA use the same learning rate but a batch size of 16 for 20 epochs. We spent 74 and 290 seconds training 1 epoch on 1,000 samples for TQA and 4,319 samples for VQA respectively. The training was performed on 1 GPU with 16GB memory of a DGX1 server ((Pascal) Tesla P100).
5 Results and Analysis
@@
In this section, we analyze the performance of our method, taking into account its distinctive characteristics as described in Table 1. Concretely, we focus on comparing our models in terms of task-specific performance, data efficiency, and multimodal capabilities.
5.1 Main Results
@@
Task-specific Performance. Table 5 and 6 demonstrate the task-specific performance of InfoSel, base models and ensemble baselines on textual and visual QA datasets. For TQA, we observe that LLaMA underperforms other base models. Upon closer examination, we found that LLaMA generates longer explanation text which, although often accurate, decreases the EM and F1-score values. Conversely, a more consistent performance of base models is observed for VQA. Mini-Viz contains 28% of unanswerable questions, and the label “unanswerable” has never been seen by base models. Consequently, this lack of exposure leads to significantly lower performance scores.
@@
Baseline ensemble methods such as WV, PageRank and OLA achieve only marginal improvements compared to base models (+1.5%) on VQA datasets. These results highlight the limitations of these methods when applied to task-specific datasets (see also Table 1). PairRanker underperformed InfoSel even though it has been trained with the same data. 0-shot LLM-Blender tends to generate longer answers compared to PairRanker which leads to a low score especially when evaluated in the exact-match settings (EM, ACC).
For TQA task, InfoSel-TT achieves 5.19% (63.63-58.44) improvement compared to individual base models, and reaches 96.12% (63.63/66.20) of the Oracle on Mini-SDv2. Similarly, the corresponding improvement in Mini-NQ performance is 1.83%, reaching 93.06% of the performance achieved by the Oracle. In contrast, FT-TT, despite its superior performance over two base models on Mini-SDv2, underperforms InfoSel-TT by more than 15% due to the small-size training data (refer to Figure 3). This low performance of FT-TT models negatively impacts the performance of InfoSel∗. As a result, we conclude that while InfoSel∗ can exhibit superior performance (see further for VQA task), it also requires more training data.
For VQA task, the results in Table 5 showcase an improvement of 4.56% in the accuracy score of InfoSel-MT compared to the base models (55.16-50.60) on Mini-GQA. Furthermore, FT-MT improves 30.48% (51.76-21.28) accuracy on Mini-Viz due to a high percentage of unseen labels (e.g., “unanswerable”) introduced during fine-tuning. Finally, the superior performance of InfoSel∗-MT model on Mini-Viz dataset demonstrates the effectiveness of the proposed blending approach, which improves 31.63% (52.91-21.28) accuracy upon the InfoSel-MT model.
@@
Data Efficiency. The experimental results shown in Figure 3 and 4 demonstrate the data efficiency of our method by evaluating the performance of the models across varying training data sizes. We observe that InfoSel-TT achieves a higher F1-score compared to the best base model when trained on as little as 10 samples from Mini-SDv2. This result has been further verified with the mean F1-score of 10 test results using different seed variations for sampling the training data (shown in Figure 6). Conversely, the number of training samples needed to surpass the performance of base models is higher for VQA datasets: 5% (6,603 samples) for Mini-GQA and 20% (864 samples) for Mini-Viz. We hypothesize that this is due to the inherent complexity of the VQA task.
@@
PairRanker is less data-efficient than InfoSel as it only achieves close performance when the training samples are more than 500 on both Mini-SDv2 and Mini-NQ. Additionally, we find that a larger training data size benefits FT-TT more than InfoSel-TT and OLA. For example, the F1-score of FT-TT increases 200% and 500% from 10 to 10,000 training samples on Mini-SDv2 and Mini-NQ respectively, while InfoSel-TT shows only marginal improvements of 3% and 4%. However, FT-TT still underperforms the best base model, which suggests that fine-tuning a small-size model requires larger training data for getting a comparable performance with LLMs or InfoSel. Finally, we observe that a larger training data size does not necessarily lead to improved performance for the fine-tuned FT-TT model (e.g., when increasing from 80% to 100% the training data size on Mini-GQA). In contrast, OLA does not benefit as much as InfoSel and FT from a larger size of training data, only outperforming InfoSel-TT on Mini-NQ when utilizing 10 and 20 training samples.
@@
Multimodal Data. InfoSel is able to utilize multimodal data (image and text) for VQA tasks, and thus outperform the latest text-exclusive LLM ensemble methods (PairRanker and LLM-Blender) as evidenced in Table 5. In contrast, PairRanker and LLM-Blender cannot process image features, thereby lacking crucial information in the multimodal setting, leading to a low accuracy on VQA datasets.777The most frequent answer of LLM-Blender on VQA datasets is “I’m sorry, I don’t have enough context to answer that question.” Further insights into the significance of modality information are elaborated in Section 5.2.
5.2 Importance of Different Input Information Fused for Ensembling
In this section, we analyze what input information is important to fuse to improve ensembling performance.
For results in Table 5, we observe that all InfoSel-TT models using different kinds of input information surpass the base models and baselines. The performance differences in using different input information are quiet margin (<0.7%). The best performance is achieved by using both the question and answer as fused input information for InfoSel-TT on Mini-SDv2. Differently, the question information works the best training on Mini-NQ.
Table 6 presents the performance results of the InfoSel-MT models utilizing various kinds of input information during the ensemble training process. Our model outperforms PairRanker when using the same kind of input information - namely, the question and answer. Additionally, incorporating the answer information alongside any supplementary data (whether visual or textual) enables our model to achieve superior performance compared to baseline models and existing benchmarks. The setting that yields the lowest accuracy solely utilizes the image (V) as the signal. This lower performance can be attributed to the fact that a single image often corresponds to multiple questions in VQA datasets, making it challenging for the model to acquire discriminative features. Furthermore, we conclude that the superior performance of our model when utilizing image, question, and answer (V+Q+A) data demonstrates the effectiveness of our model in multimodal setting.
5.3 Input Information Concatenation or Fusion
We studied the impact of concatenating and fusing multi-modal input information for the VQA task. InfoSel-MLP is an alternative model type for InfoSel which processes all the input information separately with a simple Multilayer perceptron (MLP) instead of MT. A pre-trained Sentence-BERT [32] 888https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-dot-v1 is used for generating question embedding and answer embeddings .
MLP takes the concatenated representation of question, answer, and visual embeddings as input and maps it to the label space. The objective function of InfoSel-MLP is formalized as:
The input layer of the MLP maps the concatenated representations to a hidden layer with a dimension equal to 300, followed by a ReLU activation layer and then an output layer with an output size equal to the number of models.
Table 7 presents the performance comparison between the input concatenation method (InfoSel-MLP) and the fusion method (InfoSel-MT). Our results indicate that InfoSel-MT achieves approximately 3% higher accuracy on the Mini-GQA dataset and about 2% higher accuracy on the Mini-Viz dataset compared to InfoSel-MLP. This suggests that the fused contextual representation of inputs provides more discriminative information than simply concatenating input embeddings.
Model | Input | Mini-GQA | Mini-Viz |
Fuse | ACC | ||
InfoSel-MLP | ✗ | 52.35 | 21.12 |
InfoSel-MT | 55.16 | 23.16 |